A pendulum of length 𝑙 at an angle 2𝛼. Find the time period T of the pendulum. Also let 𝛼 β†’ 0 and obtain the well-known formula 𝑇 = 2πœ‹βˆš 𝑙 /𝑔 .


1
Expert's answer
2021-12-20T11:25:27-0500

"F = ma
\\
m{a_x} = - T\sin \theta 
\\
\text{If } \sin \theta \sim \theta 
\\
m{a_x} = - T\theta 
\\
\text{Taking } \theta = \frac{x}{l}
\\
m{a_x} = - T\frac{x}{l} \qquad (i)
\\
m{a_y} = T\cos \theta - mg
\\
\text{If } {a_y} = 0, T = mg
\\
\begin{array}{l}
{a_x} = \frac{{ - gx}}{l} = - {\omega ^2}x\\
 \Rightarrow \omega = \sqrt {\frac{g}{l}} \\
T = \frac{{2\pi }}{\omega } = 2\pi \sqrt {\frac{l}{g}} 
\end{array}
\\
\text{If } \theta = 2\alpha ,T = 2\pi \sqrt {2\alpha }"